OpenStudy (anonymous):

Where do i start? Show how to apply the order of operation rules as you simplify the following expression. |-5| – 45 ÷ 3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

start by writing \[5-45\div 3\] since \(|-5|=5\)

OpenStudy (misty1212):

then divide then subtract

OpenStudy (anonymous):

45/3= 15 -5 =10?

OpenStudy (johnweldon1993):

Not quite.. \[\large 5 - 45 \div 3\] \[\large 5 - (\frac{45}{3})\] \[\large 5 - 15 = ?\]

OpenStudy (anonymous):

10?

OpenStudy (johnweldon1993):

15 is bigger than 5...so subtracting 15 from 5 would give us a "negative" number right?

OpenStudy (anonymous):

Yeah DX Forgot about that, it's 20

OpenStudy (anonymous):

-10 >~<

OpenStudy (johnweldon1993):

There we go!